If the matrix A - |--2 15 ]] i e., A= PDP with P degalizable, invertible and D = 16 d. do=d, diagonal, which choice for P?
Answer #1

1 Solution A= y 9 - 1 - 2 AI-A 1-4 g = -1 Ita (-4) ((1+2) + 9 =0 12 +21-41-8+9 =0 1²-24 + 1 =0 ta 11 - so DE 1>1 L 0 Eigen ve118 2 Ax= 3 6 (5) - lx --1 -3 - 1/3 Agein = 1 Ax= 15 (15) AX -5 Eigen Value = 1 Eigen value=1 6 6 = Eigen vector Eigen vector